OpenStudy (anonymous):

Which of the following is an equivalent form of the compound inequality −22 > −5x − 7 greater than or equal to −3? −5x − 7 < −22 and −5x − 7 greater than or equal to −3 −5x − 7 > −22 and −5x − 7 greater than or equal to −3 −5x > −22 and −7 greater than or equal to −3 −5x − 7 < −22 and −5x − 7 less than or greater to −3

OpenStudy (anonymous):

\[−22 > −5x − 7 \geq−3\] add 7 \[-15>-5x\geq 4\] divide by \(-5\) switch the inequality get \[3<x\leq-\frac{4}{5}\]